OpenStudy (anonymous):

differentiat the function y=(5-x)^72

OpenStudy (agreene):

d/dx = 72(5-x)^71

OpenStudy (phi):

Use the chain rule \[\frac{dy}{dx}= \frac{d}{dx}(5-x)^{72}=72(5-x)^{71}\frac{d}{dx}(5-x)\] \[72(5-x)^{71}\frac{d}{dx}(5-x)= 72(5-x)^{71}\cdot -1= -72(5-x)^{71}\]

OpenStudy (agreene):

yeah, should be -72(5-x)^(71) or 72(x-5)^(71) forgot my negative.